INAF LBC+MODS+LUCI queue observing, 2013 January 13-14

Observer: M. Faccini, R. Speziali, F. Cusano
Telescope Operator: D. Gonzalez Huerta
Support Astronomer: M. Edwars
Instrument: MODS and LBC

Summary

Heavy clouds on some part of the sky at sunset. We open dome at 12:40UT.
External temperature is about -16 degres, difference between primary morrors temperature and external temperature is about 8 degres
Observation started with seeing 1.8"
We close the Dome due to technical problem and low temperature -17, the dome was stucked and we close the dome by "hands"

Data Summary

Project_RUN Objects Filter ExpTime Notes
13 A3-A4 Battaglia LGS3 V and I 30'  
27D Giallongo CLJ0956+41 g and r 1h some images elongated and with bad seeing

Details

12:40 Dome Open

12:50 taking sky flats

01:32 starting observation on program 13 A3-A4 Battaglia

01:35 We slew to the target, launched pointing LGS3_pointingcheck.ob

01:37 dofpia started

01:40 dofpia does not converg, we lunch it again

01:43 still not converging we try manually with the TO, we did try with dofpia and manually until we reach the convergence at 03:29

03:30 launched LGS3_pointingcheck.ob

03:40 launched LGS3_PAm90_1.ob

03:51 stopping LGS3_PAm90_1.ob because in the red camera there was coma

03:52 launched LGS3_pointingcheck.ob

03:53 dofpia started

04:00 launched LGS3_PAm90_1.ob

04:37 Switch to MODS

05:00 switch complete we have problem with MODS, it doesn' t respond

05:31 Problem with MODS not solved, we switch to LBC

05:55 Slewing to Giallongo D

05:55 Launched pointing CLJ0956+41_GR_POINTING.ob

06:00 dofpia started

06:10 dofpia started again because the blue camera has elongated objects

06:13 Launched pointing CLJ0956+41_GR_POINTING.ob

06:17 Slewing to science, launched CLJ0956+41_GR_a.ob

07:00 dofpia started

07:04 launched CLJ0956+41_GR_b.ob

07:10 jump blue image lbcb.20130114.070518.fits

07:53 launched CLJ0956+41_GR_c.ob

07:53 dofpia started

07:56 dofpia stopped, reached the limit of x,y,z of the collimation

08:02 dofpia started

08:05 dofpia doesn't work, because in collimation we reach the limit in z. probably due to the external temperature is going down. TO is asking to software developer

08:32 launched a pointing script of Grazian just to check if the problem with dofpia is due to the telescope position

08:34 Clouds

08:50 Closing doom for clouds and low temperature, the doors of the dome are stuck temperature is -17degrees

08:51 we have to close the dome "by hands" .......... are frized

-- MarcoFaccini - 14 Jan 2013
Topic revision: r4 - 14 Jan 2013, MarcoFaccini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback